OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two.(x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (anonymous):

(x + 4)(x − 4) = x^2-16 (x + 4)(x + 4) = x^2 +8x +16 Hope that helps :)

OpenStudy (anonymous):

but how would i put that in complete sentences. And can you explain to me how to find differences and similarities so i can do it myself next time please lol :)

OpenStudy (anonymous):

@locheline also how do I give you a medal or something lol

OpenStudy (anonymous):

You click 'best response'...I'm new too, but I just hovered over it and figured it out. Conveniently XD

OpenStudy (anonymous):

Aaaand in the first one, the middle and outside terms cancel each other out because there's a negative number in there, but in the second one they are added together because both terms end up positive when multiplied.

OpenStudy (anonymous):

thanks so much lol thats awesome @locheline